Please help solve this problem:
A garden is made up of a rectangle that is 20 yards by 15 yards with a semicircle on each end.
how much fencing is needed?
I know the answer is 171.1 yards but i dont know how to get that answer please help me

Answer by mananth(16946)

let the semi circle be at the width ends
width of rectangle = 15 yards
so radius of semicircle = 7.5 yards
semicircle on either side when you add up = one circle
perimeter of circle = 2 * pi * r
=2*pi*7.5
=15pi
add to this the lengths of the rectangle
20+20=40
Total fencing required = 40 + 15pi
=87.12 yards approx.